# Independent researcher, Grenoble, France

### v1

@@
Independent researcher, Grenoble, France
@@ 1 Introduction
One of the products of academic research in static analysis is the implementation of tools, which are used to illustrate and evaluate approaches. During their lifetime, these tools have to be maintained – in particular to enable further research. This research can be performed by the same authors, new members of a same group or by other groups wanting to build upon this tool. While necessary, debugging and maintenance of static analyzers can quickly turn out to be time-consuming, as static analyzers perform highly technical reasoning. In addition, this maintenance is not the main purpose of academic jobs – and similarly, industrial developers will be looking to reduce their maintenance costs as much as possible. Due to its nature, the maintenance problem is common to all researchers in the community. Some practices are considered as folklore, which might explain why practices developed after years of experience are not systematically documented nor shared between groups. In this paper, we document the tools and techniques we use to simplify our maintenance of Mopsa, the static analysis platform we have been developing since 2017. We hope this article, following the precursor work of Andreasen et al. (2017), will have a twofold impact, by being useful to other researchers (and especially newcomers such as students), and by motivating other groups to share their practices.
@@ 2.1 Languages
Instead of an initial syntactic translation, as performed by classic analyzer frontends, Mopsa can rewrite a statement to another languages during the analysis. This dynamic translation can exploit the inferred values of variables to produce more precise, or more efficient transformations. For example, an integer addition in C can be translated to a mathematical addition if the sum of the operands fits within the range of the expression type. Similarly, an addition in Python can be translated to a mathematical addition if the operands have only values of integer types, and do not implement the special methods __add__ and __radd__. Mopsa provides ready-to-use abstractions for mathematical integers, such as intervals and polyhedra. Therefore, both C and Python analyses can delegate the processing of the translated expression to one of these domains, which enables abstraction reuse between different analyses.
@@ 3.1 Traditional Approaches to Measuring Precision
Note that other approaches to quantify precision can be leveraged when the goal is to compare the results of two analyses of a same program. One can measure and compare abstract states at similar points of the analyses. This measure however yields more questions than it solves: which abstract states should be compared? For example, comparing all abstract states in the same contexts might be time-consuming due to the sheer number of states. In addition, some precision changes will propagate from one state to its successor(s), which may skew the measure. Additional issues can arise when incomparable domains are used. In related work, the clam static analyzer from Gurfinkel and Navas (2021); Gurfinkel et al. (2015) provides a clam-diff utility to compare two different analyses of a same program, which relies on a semantic comparison of the numerical abstract states computed in each analysis. This approach has more granularity than the one we currently use, and Arceri et al. (2023) have used it to evaluate precision changes in their work. However, this approach may be too sensitive and it could be interesting to quantify the difference between some comparisons. For example, the abstract state $a_{1}=x<1$ is included in $a_{2}=x\leq 1$, which itself is included in $a_{3}=\top$, but the change from $a_{2}$ to $a_{3}$ denotes a bigger precision loss. To the best of our knowledge, this problem of quantifying the difference between two abstract states has only been considered by Sotin (2010).
@@ 4.2 Profiling
The loop hooks tracks the number of times a given loop is called, the number of iterations needed to reach a fixpoint, as well as the total time spent analyzing each loop. This is helpful to fine-tune the widening, to track slowdowns in the postfixpoint computation performed by Mopsa, and more extremely, non-termination issues.
@@ 5 Abstract Debugging through an Interactive Engine
Combined with a terminal multiplexer such as tmux, this interactive interface can be used to perform some side-by-side debugging of the analysis of a same program in two different configurations. Thanks to some wrapper scripts we have developed, commands for the interactive engine have to be entered once and will be given to both sides of the analysis, further easing debugging.
@@ 6 Automated Testcase Reduction
We start by describing two ways to leverage automated testcase reduction to ease debugging in Mopsa in Section 6.1. Our first usecase concerns internal errors from the analyzer that can happen deep into long analyses. Our second usecase pinpoints soundness issues by comparing two analyses, yielding different results on a same program. We finish by highlighting in Section 6.2 that the collaboration between Mopsa and creduce is a two-way street, as Mopsa can simplify testcase reduction for multi-file projects with complex commands such as GNU coreutils.
@@ 7 Related Work
Bugariu et al. (2018) perform automated testing of numerical abstract domains, by checking that some chosen properties should be verified. We rely on a similar yet simplified approach in Section 4.3, through an encoding of a heuristic rule to detect unsoundness during an analysis. However, our approach is not specific to numerical abstract domains.
@@ 8 Conclusion
  • Holter et al. (2024) Holter K, Hennoste JO, Lam P, Saan S, Vojdani V (2024) Abstract debuggers: Exploring program behaviors using static analysis results. In: Edwards J (ed) Proceedings of the 2024 ACM SIGPLAN International Symposium on New Ideas, New Paradigms, and Reflections on Programming and Software, Association for Computing Machinery, Onward! 2024, to appear
@@
  • Saan et al. (2024b) Saan S, Schwarz M, Erhard J, Tilscher S, Holter K, Vogler R, Apinis K, Vojdani V (2024b) Goblint analyzer documentation on testcase reduction. URL https://goblint.readthedocs.io/en/stable/developer-guide/debugging/
    #debugging-issues-with-larger-programs

### v2

# LIP6, Sorbonne Université, F-75005, Paris, France

@@
LIP6, Sorbonne Université, F-75005, Paris, France
@@ 1 Introduction
One of the products of academic research in static analysis is the implementation of tools, which are used to illustrate and evaluate approaches. During their lifetime, these tools have to be maintained – in particular to enable further research. This research can be performed by the same authors, new members of the same group or by other groups wanting to build upon this tool. While necessary, debugging and maintenance of static analyzers can quickly turn out to be time-consuming, as static analyzers perform highly technical reasoning. In addition, this maintenance is not the main purpose of academic jobs – and similarly, industrial developers will be looking to reduce their maintenance costs as much as possible. Due to its nature, the maintenance problem is common to all researchers in the community. Some practices are considered as folklore, which might explain why practices developed after years of experience are not systematically documented nor shared between groups. In this paper, we document the tools and techniques we use to simplify our maintenance of Mopsa, the static analysis platform we have been developing since 2017. We hope this article, following the precursor work of Andreasen et al. (2017), will have a twofold impact, by being useful to other researchers (and especially newcomers such as students), and by motivating other groups to share their practices.
@@ 2.1 Languages
Instead of an initial syntactic translation, as performed by classic analyzer frontends, Mopsa can rewrite a statement to other languages during the analysis. This dynamic translation can exploit the inferred values of variables to produce more precise, or more efficient transformations. For example, an integer addition in C can be translated to a mathematical addition if the sum of the operands fits within the range of the expression type. Similarly, an addition in Python can be translated to a mathematical addition if the operands have only values of integer types, and do not implement the special methods __add__ and __radd__. Mopsa provides ready-to-use abstractions for mathematical integers, such as intervals and polyhedra. Therefore, both C and Python analyses can delegate the processing of the translated expression to one of these domains, which enables abstraction reuse between different analyses.
@@ 3.1 Traditional Approaches to Measuring Precision
Note that other approaches to quantify precision can be leveraged when the goal is to compare the results of two analyses of the same program. One can measure and compare abstract states at similar points of the analyses. This measure however yields more questions than it solves: which abstract states should be compared? For example, comparing all abstract states in the same contexts might be time-consuming due to the sheer number of states. In addition, some precision changes will propagate from one state to its successor(s), which may skew the measure. Additional issues can arise when incomparable domains are used. In related work, the clam static analyzer from Gurfinkel and Navas (2021); Gurfinkel et al. (2015) provides a clam-diff utility to compare two different analyses of the same program, which relies on a semantic comparison of the numerical abstract states computed in each analysis. This approach has more granularity than the one we currently use, and Arceri et al. (2023) have used it to evaluate precision changes in their work. However, this approach may be too sensitive and it could be interesting to quantify the difference between some comparisons. For example, the abstract state $a_{1}=x<1$ is included in $a_{2}=x\leq 1$, which itself is included in $a_{3}=\top$, but the change from $a_{2}$ to $a_{3}$ denotes a bigger precision loss. To the best of our knowledge, this problem of quantifying the difference between two abstract states has only been considered by Sotin (2010).
@@ 3.2 Reporting static analysis results in an automatic and transparent fashion
When counting the number of checks (total and proved safe), it is possible to count the same location several times, as our analysis is fully context-sensitive: every context (i.e., call-stack) is counted separately. In general, comparing only the sensitivity of analyses with different coverage, causing the number of contexts to vary, can yield meaningless results. This justifies the introduction of fine-grained result comparison in the next subsection for those cases. However, although Mopsa performs loop unrolling to improve precision, this does not affect the selectivity as unrolling is not part of the context: a check is considered safe if it is safe for all unrollings of the same context.
@@ 4.2 Profiling
The loop hook tracks the number of times a given loop is called, the number of iterations needed to reach a fixpoint, as well as the total time spent analyzing each loop. This is helpful to fine-tune the widening, to track slowdowns in the postfixpoint computation performed by Mopsa, and more extremely, non-termination issues.
@@ 5 Abstract Debugging through an Interactive Engine
Combined with a terminal multiplexer such as tmux, this interactive interface can be used to perform some side-by-side debugging of the analysis of the same program in two different configurations. Thanks to some wrapper scripts we have developed, commands for the interactive engine have to be entered once and will be given to both sides of the analysis, further easing debugging.
@@ 6 Automated Testcase Reduction
We start by describing two ways to leverage automated testcase reduction to ease debugging in Mopsa in Section 6.1. Our first usecase concerns internal errors from the analyzer that can happen deep into long analyses. Our second usecase pinpoints soundness issues by comparing two analyses, yielding different results on the same program. We finish by highlighting in Section 6.2 that the collaboration between Mopsa and creduce is a two-way street, as Mopsa can simplify testcase reduction for multi-file projects with complex commands such as GNU coreutils.
@@ 7 Related Work
Bugariu et al. (2018) perform automated testing of numerical abstract domains, by checking that some chosen properties should be verified. We rely on a similar yet simplified approach in Section 4.3, through an encoding of a heuristic rule to detect unsoundness during an analysis. However, our approach is not specific to numerical abstract domains. Goblint relies on a similar approach implemented by Saan and Schwarz (2022) to check that at least one branch in conditionals is not dead.
@@ 8 Conclusion
  • Holter et al. (2024) Holter K, Hennoste JO, Lam P, Saan S, Vojdani V (2024) Abstract debuggers: Exploring program behaviors using static analysis results. Association for Computing Machinery, Onward! ’24, p 130–146, DOI: 10.1145/3689492.3690053
@@
  • Saan and Schwarz (2022) Saan S, Schwarz M (2022) URL https://github.com/goblint/analyzer/issues/826
@@
  • Saan et al. (2024b) Saan S, Schwarz M, Erhard J, Tilscher S, Holter K, Vogler R, Apinis K, Vojdani V (2024b) Goblint analyzer documentation on testcase reduction. URL https://goblint.readthedocs.io/en/stable/developer-guide/debugging/#debugging-issues-with-larger-programs